Find the determinant of each matrix. i cant find how to do this in my book. so please write the steps with answers. thank you


[-6 -7]
[5 8]


[1 -5]
[-2 9]


[1 2 3]
[6 0 5]
[0 4 0]

Answer by stanbon(75887)

Use Google with search string "matrix determinant"
to see how determinants are calculated.
-------------
Your Problems:
[-6 -7]
[5 8]
D = (-6*8)-(-7*5) = -48 + 35 = -13
-----------------------------------------
[1 -5]
[-2 9]
D = (1*9)-(-5*-2) = 9 - 10 = -1
------------------------------------
[1 2 3]
[6 0 5]
[0 4 0]
D = (1*0*0 + 6*4*3 + 0*5*2)-(3*0*0 + 5*4*1 + 0*6*2)
= (0+72+0) - (0+20+0) = 52
-----------------------------------
Cheers,
Stan H.